### 13 Mar 2024 v1

# Language models scale reliably with over-training and on downstream tasks

Abstract
Scaling laws are useful guides for developing language models, but there are still gaps between current scaling studies and how language models are ultimately trained and evaluated. For instance, scaling is usually studied in the compute-optimal training regime (i.e., “Chinchilla optimal” regime); however, in practice, models are often over-trained to reduce inference costs. Moreover, scaling laws mostly predict loss on next-token prediction, but ultimately models are compared based on downstream task performance. In this paper, we address both shortcomings. To do so, we create a testbed of 104 models with 0.011B to 6.9B parameters trained with various numbers of tokens on three data distributions. First, we investigate scaling in the over-trained regime. We fit scaling laws that extrapolate in both the number of model parameters and the ratio of training tokens to parameters. This enables us to predict the validation loss of a 1.4B parameter, 900B token run (i.e., 32$\times$ over-trained) and a 6.9B parameter, 138B token run—each from experiments that take 300$\times$ less compute. Second, we relate the perplexity of a language model to its downstream task performance via a power law. We use this law to predict top-1 error averaged over downstream tasks for the two aforementioned models using experiments that take 20$\times$ less compute. Our experiments are available at https://github.com/mlfoundations/scaling.
00footnotetext: Equal advising, ordered alphabetically. Correspondence to sy@cs.columbia.edu. ${}^{1}$olumbia University ${}^{2}$oyota Research Insitute ${}^{3}$ Austin ${}^{4}$pple ${}^{5}$niversity of Washington ${}^{6}$uelich Supercomputing Center, Research Center Juelich ${}^{7}$ ${}^{8}$tanford University ${}^{9}$el Aviv University ${}^{10}$ Munich ${}^{11}$ontextual AI
1 Introduction
Figure 1: Reliable scaling in the over-trained regime and for downstream error prediction. (left) We fit a scaling law for model validation loss, parameterized by (i) a token multiplier $M$, which is the ratio of training tokens $D$ to parameters $N$ (i.e., $M=D/N$) and (ii) the compute $C$ in FLOPs used to train a model, approximated by $C=6ND$. We extrapolate, in both $N$ and $M$, the validation performance of models requiring over 300$\times$ the training compute used to construct the scaling law. (right) We also fit a scaling law to predict average downstream top-1 error as a function of validation loss. We find that fitting scaling laws for downstream error benefits from using more expensive models when compared to fitting for loss prediction. We predict the average error over 17 downstream tasks for models trained with over 20$\times$ the compute.
Training large language models is expensive. Moreover, training high-quality models requires a complex recipe of algorithmic techniques and training data. To reduce the cost of finding successful training recipes, researchers first evaluate ideas with small experiments and then extrapolate their efficacy to larger scales. With reliable extrapolation, it is possible to quickly iterate at small scale and still pick the method that will perform best for the final large training run. Indeed, this workflow has become commonplace for training state-of-the-art language models such as Chinchilla 70B [43], PaLM 540B [17], and GPT-4 [74].
Despite their importance for model development, published scaling laws differ from the goals of training state-of-the-art models in important ways. For instance, scaling studies usually focus on the compute-optimal training regime (“Chinchilla optimality” [43]), while widely used models are now often over-trained to reduce inference costs. Another potential mismatch between scaling laws and eventual applications of the models is that most scaling laws quantify model performance by perplexity in next-token prediction instead of accuracy on widely used benchmark datasets. As a result, it is unclear whether following scaling laws leads to truly better models, or merely to models with lower perplexity in the compute-optimal training regime.
In this paper, we address both topics: scaling in the over-trained regime and downstream performance prediction, with an extensive set of experiments.
Motivated by the practice of training beyond compute-optimal, we first investigate whether scaling follows reliable trends in the over-trained regime. We find that for a set of model configurations with a constant ratio of training tokens to parameters, the models’ reducible loss $L^{\prime}$ [41, 43] follows consistent power laws ($L^{\prime}=\lambda\cdot C^{-\alpha_{C}}$) in the amount of training compute $C$. As one increases the ratio of tokens to parameters, corresponding to more over-training, the scaling exponent $\alpha_{C}$ remains about the same, while the scalar $\lambda$ changes. We consider the extent to which our observations are explainable and find a promising approach by reparameterizing forms for scaling laws.
To establish if and when scaling is predictable in the over-trained regime, we experiment with a testbed of 104 models, trained from scratch on three different datasets: RedPajama [110], C4 [86, 25], and RefinedWeb [80]. We find that scaling laws fit on small models, trained closer to compute-optimal, can accurately predict the performance of larger models that undergo more over-training. Figure 1 (left) illustrates our main over-training result, where we invest $192.4e$ 19 FLOPs to extrapolate the C4 validation performance of a 1.4B parameter model trained on 900B tokens, which requires 300×300$\times$300 × more compute to train.
In addition to over-training, we also investigate if scaling laws can predict the performance of a model on downstream tasks. We establish a power law relationship between language modeling perplexity and the average top-1 error on a suite of downstream tasks. While it can be difficult to predict the error on individual tasks, we find that aggregate performance can be accurately predicted from a model’s perplexity among models trained on the same training data. Figure 1 (right) presents our main downstream error prediction result, where we invest $202.7e$ 20 FLOPs to predict the average top-1 error over a set of downstream tasks to within 1 percentage point for a 6.9B compute-optimal model, which requires 20×20$\times$20 × more compute to train.
To facilitate further research on reliable scaling, we provide all results of our experiments at https://github.com/mlfoundations/scaling.
2 Scaling and over-training
In this section, we describe empirical observations and their potential mathematical descriptions. First, we provide key definitions (Section 2.1). We next present a phenomenon wherein training a collection of models for increasing token multipliers—ratios of tokens to parameters—follows similar scaling trends (Section 2.2). We then show that these patterns are consistent with previously proposed power-law scaling when reparameterizing in terms of training compute and token multipliers (Section 2.3). Towards connecting loss scaling and downstream performance, we revisit our collection of models, plot their average top-1 error vs. validation loss, and notice that error decays exponentially with lower loss (Section 2.4). To describe these observations, we propose a scaling law for downstream error as a function of loss (Section 2.5).
2.1 Preliminaries
Scaling laws for loss.
We examine scaling laws that predict the loss $L$ of a model as a function of the compute $C$ in FLOPs used to train the model. If one increases the number of parameters $N$ in a model or the number of tokens $D$ that a model is trained on, compute requirements naturally increase. Hence, $C$ is assumed to be a function of $N,D$. Following Kaplan et al. [49], we use the approximation $C=6ND$, which Hoffmann et al. [43] independently verify. We consider scaling laws,
$L(C)=E+L^{\prime}(C),$ (1)
where $E$ is an irreducible loss and $L^{\prime}$ is the reducible loss. $E$ captures the Bayes error or minimum possible loss achievable on the validation domain. The $L^{\prime}(C)$ term captures what can possibly be learned about the validation domain by training on a source domain. $L^{\prime}(C)$ should go to zero with increased training data and model capacity. $L^{\prime}(C)$ is often assumed to follow a power law: $L^{\prime}(C)=\lambda\cdot C^{-\alpha_{C}}$ (i.a.,  Hestness et al. [41], OpenAI [74]). It is also often helpful to consider a power law in a $\log$-$\log$ plot, where it appears as a line with slope $-\alpha_{C}$ and $y$-intercept $\log{(\lambda)}$.
Token multipliers.
We define a token multiplier $M=D/N$ as the ratio of training tokens to model parameters. We introduce $M$ for notational convenience as it allows us to consider fixed relationships between $D$ and $N$ even as a model gets bigger (i.e., as $N$ becomes larger).
Compute-optimal training.
Hoffmann et al. [43] establish compute-optimal training, where, for any compute budget $H$, the allocation of parameters and tokens that minimize training or validation loss is given by,
$\arg\min_{N,D}L(N,D)\text{ s.t. }C(N,D)=H.$ (2)
To solve for the optimal $N^{*},D^{*}$, one can sweep $N,D$ for each $H$, retaining the best configurations. Hoffmann et al. [43] find that as $H$ increases, $N^{*}$ and $D^{*}$ scale roughly evenly. Assuming equal scaling, there is a fixed compute-optimal token multiplier $M^{*}=D^{*}/N^{*}$ per training distribution.
Over-training.
We define over-training as the practice of allocating compute sub-optimally, so smaller models train on a disproportionately large number of tokens (i.e., $M>M^{*}$). While loss should be higher than in the compute-optimal allocation for a given training budget, the resulting models have fewer parameters and thus are cheaper at inference.
Figure 2: Scaling in the over-trained regime follows consistent power law exponents. We notice parallel lines in the $\log$-$\log$ plots of reducible loss vs. training compute for a range of token multipliers $M$, which give the ratio of training tokens to model parameters and where larger $M$ corresponds to more over-training. For a power law giving reducible loss as a function of compute: $L^{\prime}(C)=\lambda\cdot C^{-\alpha_{C}}$, the scaling exponent $\alpha_{C}$ remains relatively constant resulting in lines with approximately fixed slope. The $y$-intercept, which is determined by the power law coefficient $\lambda$, however, shifts with different values $M$. This suggests that the power law coefficient $\lambda$ is a function of the token multiplier $M$, while the power law exponent $\alpha_{C}$ is not. Hestness et al. [41] report a similar phenomenon of consistent scaling exponents for recurrent networks when modifying architectures instead of token multipliers.
2.2 Observation: Over-trained models follow consistent trends
We begin our scaling investigation by training models with 0.011B to 0.411B parameters for token multipliers $M$ between 20 and 640, where $M=20$ points lie roughly on the compute-optimal frontier, and $M>20$ corresponds to over-training. We defer experimental details to Section 3 to focus on our observations. In Figure 2, we plot loss against compute on a $\log$-$\log$ scale for the models trained on three datasets and evaluated on the C4 eval set. We notice a phenomenon of “parallel lines” when fitting power laws to the reducible loss, which suggests a near-constant scaling exponent even with increased over-training. This indicates that scaling in the over-trained regime may be predictable given training runs closer to compute-optimal.
2.3 Deriving scaling laws for over-trained behavior
In search of an analytic expression for the observations in Figure 2, we turn to the scaling literature. A common functional form for the risk of a model, as proposed in prior work [91, 43] is,
$L(N,D)=E+AN^{-\alpha}+BD^{-\beta}.$ (3)
Recall from Section 2.1, $N$ is the number of parameters, $D$ the number of training tokens and $E$ the irreducible loss. The constants $E,A,\alpha,B,\beta$ are fit from data. By fitting this parametric form, Hoffmann et al. [43] find that scaling exponents $\alpha$ and $\beta$ are close, suggesting that one should scale $N,D$ equally as compute increases. Hence, we assume $\alpha=\beta$. With this assumption, we reparameterize Equation (3) in terms of compute $C=6ND$ and a token multiplier $M=D/N$. We get the following form,
$L(C,M)=E+\left(aM^{\alpha_{C}}+bM^{-\alpha_{C}}\right)C^{-\alpha_{C}},$ (4)
where $\alpha_{C}=\alpha/2$, $a=A(1/6)^{-\alpha_{C}}$, $b=B(1/6)^{-\alpha_{C}}$ gives the relation to Equation (3). For a complete derivation, see Appendix B.
Equation (4) has the following interpretation: (i) The scaling exponent $\alpha_{C}$ is not dependent on $M$. Thus, we always expect lines with the same slope in the $\log$-$\log$ plot—as in Figure 2. (ii) The term $aM^{\alpha_{C}}+bM^{-\alpha_{C}}$ determines the offsets between curves with different token multipliers. Hence, we expect non-overlapping, parallel lines in the $\log$-$\log$ plot for the range of $M$ we consider—also consistent with Figure 2.
Recall that we make the assumption $\alpha=\beta$, which implies equal scaling of parameters and tokens as more compute is available. However, as explained in Appendix B, even if $\alpha\neq\beta$, we get a parameterization that implies the power-law exponent remains constant with over-training.
2.4 Observation: Loss tracks average top-1 error
Figure 3: Average top-1 error scales as a function of loss. We plot data points trained on three datasets and notice an exponential decay of average top-1 error as C4 eval loss, on the x-axis, decreases. The specific coefficients appear dataset-dependent. We consider on the y-axes (left) the average error over 46 evaluations and (right) the average error on a subset of 17 evaluations where performance can be 10 points above random chance for at least one 0.154B scale model. These observations suggest that average top-1 error should be predictable with reliable loss estimates.
Scaling is typically studied in the context of loss [49, 43, 70], which Schaeffer et al. [98] note is smoother than metrics like accuracy. However, practitioners ultimately care about downstream, in-the-wild task performance. To better connect scaling laws and over-training to task prediction, we revisit the suite of models plotted in Figure 2. In Figure 3, we plot average downstream top-1 errors over evaluations sourced from LLM-Foundry [67] against the C4 eval loss. We defer details of the setup to Section 3 to focus here on a few key observations. The average errors appear to follow exponential decay as loss decreases. Additionally, the particular relationship between loss and error is dataset-dependent. For instance, models trained on C4 result in the lowest C4 eval loss, but this does not translate to downstream gains compared to models trained on RedPajama or RefinedWeb.
2.5 Proposing a scaling law for average top-1 error
Based on the exponential decay we observe in Figure 3, we propose the following relationship between downstream average top-1 error Err and loss $L$,
$\textsf{Err}(L)=\epsilon-k\cdot\exp{(-\gamma L)},$ (5)
where $\epsilon,k,\gamma$ are fit from data. Equation (5) also has an appealing interpretation in terms of model perplexity $PP(L)=\exp{(L})$,
$\textsf{Err}(PP)=\epsilon-k\cdot PP^{-\gamma}.$ (6)
Namely, Err follows a power law in $PP$ with maximum error $\epsilon$, where intuitively $\epsilon$ should be close to the random chance performance.
Equation (5) in conjunction with (4) suggests a two-step method to predict Err as a function of compute and the amount of over-training. For choices of training and validation distributions, (i) fit a scaling law to Equation (4) using $((C,M),L)$ pairs to yield $(C,M)\mapsto L$. (ii) Fit a scaling law to Equation (5) using $(L,\textsf{Err})$ pairs to get $L\mapsto\textsf{Err}$.
3 Experimental setup
Towards testing the analytic predictions in Equations (4), (5), we discuss our experimental setup. We first present language modeling details (Section 3.1). Next, we discuss our strategy for deciding which models to include in our scaling investigation and our procedure for fitting scaling trends (Section 3.2). We then present metrics to validate how well scaling laws predict loss and downstream performance (Section 3.3).
3.1 Training setup
Architecture.
We train transformers [114], based on auto-regressive, decoder-only, pre-normalization architectures like GPT-2 [83] and LLaMA [111]. We adopt OpenLM [37] as our core modeling library, which utilizes PyTorch [78, 6], xformers [52], triton [73], FlashAttention [22], FSDP [127], and bfloat16 automatic mixed precision. Like LLaMA, we omit bias terms, but replace RMSNorm [125] with LayerNorm [8], which has readily available fused implementations. Following Wortsman et al. [120], we apply qk-LayerNorm [23], which adds robustness to otherwise poor hyperparameter choices (e.g., for learning rate). We use SwiGLU [100] activations and depth-scaled initialization [126]. We use a sequence length of 2048, rotary positional embeddings [104], and the GPT-NeoX-20B tokenizer [13], which yields a vocabulary size of 50k. We do not use weight tying [82, 44].
Objectives and optimization.
[…]
Training datasets.
To ensure our conclusions are not particular to a training distribution, we train models on C4 [86, 25], RedPajama [110], and RefinedWeb [80]. They are open-source and have 138B, 1.15T, and 600B tokens respectively. We sample without replacement and employ sequence packing without attention masking. We separate documents in our training corpora with end-of-text tokens.
3.2 Creating scaling laws for validation loss and downstream error prediction
Figure 4: Search, filter, fit: A recipe for selecting configurations for scaling. (left) To generate the final configurations presented in Table 1, we run a 435 model grid search over model width, hidden dimension, number of attention heads, batch size, and warmup steps. All models are trained near compute-optimally. (center) We plot the efficient frontier of models, which appear to follow a trend, excluding models from $5.2\times 10^{16}$ to $5.2\times 10^{17}$, which fall below the trend. (right) We fit a power law with irreducible error to the remaining configurations, picking four configurations that closely track the full model suite. These models extrapolate the performance of 1.4B, 6.9B target models. Shaded regions represent bootstrap 95% confidence intervals.
Recipe for model configuration selection.
To create a testbed of models for our scaling experiments, we grid search over a total of 435 models, trained from scratch, in the 0.01B to 0.5B parameter range as seen in Figure 4 (left). We train on the OpenLM data mix [37], which largely consists of tokens from RedPajama [110] and The Pile [29]. We train on 20 tokens per parameter ($M=20$), which we find in early experiments gives models near the compute-optimal frontier for the data mix. This is similar to findings presented in Hoffmann et al. [43]’s Table 3, which suggests that roughly 20 tokens per parameter are optimal in their experimental setup. Our validation set, OpenLM eval, contains tokens from recent arXiv papers, the OpenLM codebase itself, and news articles. To find maximally performant models on validation data, we tune model width, number of layers, number of attention heads, warmup steps, and batch size. We find in early experiments that qk-LayerNorm makes models less sensitive to learning rate, which is a phenomenon Wortsman et al. [120] report in their Figure 1. Hence, we fix the learning rate for our sweeps. We also perform smaller grid searches over 1.4B and 6.9B parameter model configurations at $M=20$, retaining the best configurations.
In Figure 4 (center), we plot the efficient frontier of minimum loss configurations. While there appears to be a general trend, configurations between $5.2\times 10^{16}$ and $5.2\times 10^{17}$ FLOPs lie below the frontier established by other models. We hypothesize these models over-perform as they are trained for more optimization steps than their neighbors based on our power-of-two batch sizes. We provide support for this hypothesis in Appendix E, but opt to remove these models from our investigation.
In Figure 4 (right), we fit trends to the remaining models and to a subset of four models. We notice that the trends hit both the 1.4B and 6.9B models suggesting that our small-scale configurations are reasonable for extrapolation to larger parameter and compute regimes. We retain the four model configuration subsets as a representative sample.
We do not tune hyperparameters for other token multipliers (i.e., $M\neq 20$), other training or evaluation distributions, or on downstream task validation sets. For more details, see Appendix C. We present our final hyperparameters in Table 1 given their importance.
$N$ $n_{\text{layers}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{head}}$ Warmup Learning rate Batch size $M=20$ A100 hours
[…]
Table 1: Main models and hyperparameters used in our investigation. Models have number of parameters $N$, with number of layers $n_{\text{layers}}$, number of attention heads $n_{\text{heads}}$, model width $d_{\text{model}}$, and width per attention head $d_{\text{head}}$. Batch sizes are global and in units of sequences. Each sequence has 2048 tokens. A100 GPU hours are at $M=20$, which are near compute-optimal runs. For the 1.4B scale, a batch size of 256 performs slightly better than 512.
Fitting scaling laws.
$N$ $M$ Used to fit Equation (4) Used to fit Equation (5)
0.011B 20
0.079B 20
0.154B 20
0.411B 20
0.011B 320
1.4B 20
[…]
Table 2: Default $N,M$ to fit our scaling laws. We invest $\sim$100 A100 hours to fit Equation (4) and $\sim$1000 A100 hours to fit Equation (5).
We fit Equation (4) to approximate $E,a,b,\alpha_{C}$ using curve-fitting in SciPy [115] (i.e., Levenberg-Marquardt to minimize non-linear least squares). We try several initializations and retain the best fit. We repeat this process to fit Equation (5) to approximate $\epsilon,k,\gamma$. Unless otherwise specified, we fit to the $N,M$ pairs in Table 2. In total, we invest $\sim$100 A100 hours to train the models required for fitting an accurate scaling law for loss prediction and $\sim$1000 A100 hours for a corresponding scaling law for downstream error prediction. Our configurations allow us to test for extrapolation to the $N=1.4\text{B},M=640$ (900B token) and the $N=6.9\text{B},M=20$ (138B token) regimes.
3.3 Evaluation setup
Evaluation datasets.
Unless otherwise stated, C4 eval is our default validation loss dataset. For downstream tasks, we adopt 46 tasks from LLM-foundry [67], which includes standard tasks with both zero-shot and few-shot evaluations. We also consider a 17-task subset where, for each evaluation, at least one 0.154B scale model—trained with as many as 99B tokens—gets 10 percentage points above chance accuracy: ARC-Easy [21], BIG-bench: CS algorithms [11], BIG-bench: Dyck languages [11], BIG-bench: Novel Concepts [11], BIG-bench: Operators [11], BIG-bench: QA WikiData [11], BoolQ [19], Commonsense QA [105], COPA [90], CoQA [89], HellaSwag (zero-shot) [123], HellaSwag (10-shot) [123], LAMBADA [75], PIQA [12], PubMed QA Labeled [48], SQuAD [88], and WinoGrand [53]. This subset allows us to investigate a regime where performance for small models can be non-trivial. For more details on evaluation datasets see Appendix D. For ablations on our choices of loss and downstream evaluations see Appendix E.
Metrics.
We consider three main metrics: (i) Validation loss, which is the cross entropy between a model’s output and the one-hot ground truth, averaged over all tokens in a sequence and over all sequences in a dataset. (ii) Average top-1 error, which is a uniform average over 46 downstream evaluations sourced from LLM-foundry [67]. We also look at the mean top-1 error for the subset of 17 evaluations identified in the paragraph above. For a complete list of downstream evaluation datasets, see Appendix D. To measure how good a prediction $\zeta(C,M)$ is, we measure (iii) Relative prediction error: $|\zeta(C,M)-\zeta_{GT}|/\zeta_{GT}$, where $\zeta$ is the loss $L$ or the average top-1 error Err.
Testbed.
We train models on C4, RedPajama, and RefinedWeb with the number of parameters $N\in\{0.011\text{B},0.079\text{B},0.154\text{B},0.411\text{B}\}$ and token multipliers $M\in\{5,10,20,40,80,160,320,640\}$. We omit runs that require more tokens than are present in a dataset (i.e., $N=0.411\text{B},M=640$ for C4). We additionally train $N=1.4$B models at $M=20$ and at the largest token multiplier possible without repeating tokens (i.e., 80 for C4, 640 for RedPajama, and 320 for RefinedWeb). We train $N=6.9\text{B},M=20$ for each dataset. In total this results in 104 models. We evaluate each model on C4 eval for validation loss and on the 46 downstream tasks for top-1 error.
4 Results
Unless otherwise stated, we fit Equations (4), (5) to the configurations in Table 2 and use C4 eval for loss computation.
Over-trained performance is predictable.
We highlight our main over-training results in Figure 1 (left). Namely, we are able to extrapolate both in the number of parameters $N$ and the token multiplier $M$ to closely predict the C4 eval performance of a 1.4B parameter model trained on 900B RedPajama tokens ($N=1.4\text{B},M=640$). Our prediction, which takes 300$\times$ less compute to construct than the final 1.4B run, is accurate to within 0.7% relative error. Additionally, for the $N=6.9\text{B},M=20$ run, near compute-optimal, the relative error is also 0.7%.
These results support several key takeaways. (i) scaling can be predictable even when one increases the model size and the amount of over-training compared to the training runs used to fit a scaling law. (ii) Scaling can be predictable even in the presence of a distribution shift (e.g., RedPajama training and C4 evaluation). (iii) The form presented in Equation (4) is useful in practice for fitting and predicting over-trained scaling behavior. (iv) Fitting to Equation (4) does not sacrifice prediction accuracy near compute-optimal.
Figure 5: Relative error on C4 eval for different training distributions. Boxes highlighted in yellow correspond to data points used to fit Equation (4), all other points are predicted based on the fit. The prediction error is low in both interpolation and extrapolation ranges for number of parameters $N$ and token multiplier $M$, with larger values of $M$ specifying more over-training. Empty squares correspond to runs that were not possible due to compute limitations at the 1.4B and 6.9B scales or to limited dataset size for single epoch training.
While Figure 1 explores a specific case of making predictions in the over-trained regime, we would like to understand the error profile of our predictions across datasets, token multipliers, and number of parameters. Hence, in Figure 5 we show relative error between ground truth loss and predicted loss on C4 eval for models in our testbed. We notice uniformly low prediction error suggesting that predictions are accurate in many settings.
Average top-1 error is predictable.
Figure 1 (right) presents our results in estimating scaling laws for downstream error. Similar to Figure 1 (left), we are able to extrapolate in $N,M$ and predict the average downstream error across our evaluation suite. Concretely, we use the models indicated in Table 2 to fit Equation (5), and predict the average top-1 error over the 17 tasks identified in Section 3.3. Our fit allows us to predict the downstream performance of a 6.9B parameter model trained on 138B tokens to within $0.05\%$ relative error and of a 1.4B model trained on 900B tokens to within $3.6\%$ relative error, using 20×20$\times$20 × less compute.
Table 3 additionally shows the relative error of our downstream performance predictions for models trained on C4, RedPajama, and RefinedWeb, indicating that our scaling law fits are applicable to other datasets. We note that while average accuracy across benchmarks is predictable, predicting accuracy on individual downstream tasks is significantly more noisy. We report relative error of all our predictions in Figures 11, 12 in the Appendix. We also find that if we remove the 1.4B model for the Equation (5) fit, relative error jumps, for instance, from 0.05% to 10.64% on the 17-task split for the 6.9B, 138B token RedPajama prediction.
Individual top-1 error Average top-1 error
Train set MMLU [38] OpenBook QA [66] HellaSwag [123] 17-task split 46-task split
C4 [86, 25] 2.82% 16.80% 79.58% 0.42% 0.76%
RedPajama [110] 0.12% 8.44% 25.73% 0.05% 2.10%
RefinedWeb [80] 0.77% 1.92% 81.96% 2.94% 2.76%
Table 3: Downstream relative prediction error at 6.9B, 138B tokens. While predicting accuracy on individual zero-shot downstream evaluations can be challenging (“Individual”), predicting averages across downstream datasets is accurate (“Average”). On the right, we report relative prediction error for the average over a 17-task subset and over the full 46-task suite.
Small-scale experiments can predict model rank order.
We expect to be able to rank models based on their predicted performance, which is useful when deciding what to train. To verify, we rank 9 testbed models with $N\geq 1.4\text{B}$ by ground-truth top-1 error and by estimated top-1 error. We find high rank correlations: 0.93 and 0.88 for the 46 and 17-task splits respectively.
Under-training, out-of-distribution scaling, and compute-reliability trade-offs.
In addition to our main results presented above, we include additional results in Appendix E, which we summarize here. First, we notice that when token multipliers become too small (i.e., $M=5$) scaling becomes unreliable and lies off the trend. Additionally, several multipliers (10, 20, 40, and 80) garner points that are roughly on the compute optimal frontier (Figure 9). To probe the limits of reliable scaling, we attempt to break our scaling laws in out-of-distribution settings. We find that models trained on C4—English filtered—and evaluated on next token prediction on code domains have a high relative error in many cases. Perhaps surprisingly, evaluating the same models on German next token prediction again gives reliable loss scaling (Figure 10). We additionally examine the compute necessary to create accurate scaling laws, finding a positive correlation between investing more compute in a scaling law and its predictivity. We find that scaling laws can be constructed more cheaply for loss prediction than for downstream error prediction (Figures 1516).
5 Related work
We review the most closely related work in this section. For additional related work, see Appendix F.
Model family Parameters $N$ Training tokens $D$ Token multiplier $M$
T5 [87] 11B 34B 3.1
GPT-3 [14] 175B 300B 1.7
Gopher [84] 280B 300B 1.1
Chinchilla [43] 70B 1.4T 20.0
LLaMA [111] 7B 1T 140.0
LLaMA [111] 70B 1.4T 20.0
LLaMA-2 [112] 7B 2T 290.0
LLaMA-2 [112] 70B 2T 30.0
XGen [72] 7B 1.5T 210.0
MPT [108] 7B 1T 140.0
Table 4: Token multipliers of existing models. In our work, we run experiments with token multipliers between 5 and 640 for (GPT-2 [83], LLaMA [111])-style decoder-only architectures.
Scaling laws.
Early works on scaling artificial neural networks observe predictable power-law scaling in the training set size and number of model parameters [41, 42, 91]. Alabdulmohsin et al. [2] center the importance of looking at the extrapolation regime of a scaling law. Yang et al. [121] prescribe architectural and hyperparameter changes when scaling model width to realize performant models; Yang et al. [122] make analogous recommendations when scaling model depth. Unlike the aforementioned work, our investigation focuses on the link between over-training models and predicting their downstream performance on accuracy metrics.
Hoffmann et al. [43] investigate how the number of model parameters $N$ and training tokens $D$ should be chosen to minimize loss $L$ given a compute budget $C$. Hoffmann et al. [43] find that when scaling up $C$, both $N$ and $D$ should be scaled equally up to a multiplicative constant (i.e., $N\propto C^{\sim 0.5}$ and $D\propto C^{\sim 0.5}$) to realize compute-optimality. Appendix C of the Chinchilla paper additionally suggests that these findings hold across many datasets. However, Hoffmann et al. [43] do not account for inference costs, provide scaling laws for training beyond compute-optimal, or for downstream error prediction—all of which are central to our work.
Sardana & Frankle [96] proposed modifications to the Chinchilla formulation to incorporate inference costs into the definition of compute-optimal and solve for various fixed inference budgets. Their key finding, which is critical for our work, is that when taking into account a large enough inference budget, it is optimal to train smaller models for longer when compared to the original Chinchilla recommendations. Our work presupposes that over-training can be beneficial. Instead of solving for inference-optimal schemes, we support empirically a predictive theory of scaling in the over-trained regime. Additionally, we provide experiments across many validation and training sets.
For predicting downstream scaling beyond loss, Isik et al. [45] relate the number of pre-training tokens to downstream cross-entropy and machine translation BLEU score [76] after fine-tuning. In contrast, we do not examine a specific domain but rather take a holistic approach to evaluation by looking at top-1 error over many natural language tasks. Schaeffer et al. [98] argue that emergent abilities [118] and unreliable scaling are a product of non-linear metrics like error/accuracy and propose smoother alternatives. As an explanation for why non-linear metrics may be hard to predict, Schaeffer et al. [98] consider predicting exactly an $\ell$ length sequence: $\textsf{Err}(N,\ell)\approx 1-PP(N)^{-\ell}$, where $N$ is the number of parameters in a model and $PP$ is its perplexity. This is a special case of our Equations (5), (6), where number of training tokens is ignored, $\epsilon=1,k=1$, and $\gamma=\ell$. In contrast, we treat $\epsilon,k,\gamma$ as free parameters for a scaling law fit and embrace top-1 error, finding that average error over downstream tasks can make for a predictable metric.
Over-training in popular models.
There has been a rise in over-trained models [111, 112] and accompanying massive datasets [110, 80, 102, 3]. To contextualize the extent to which we over-train, we provide token multipliers for popular models in Table 4. For example, Chinchilla 70B [43] is trained with a token multiplier of 20, while LLaMA-2 7B [112] uses a token multiplier of 290. In our investigation, we look at token multipliers from 5 to 640 to ensure coverage of popular models and relevance for future models that may be trained on even more tokens.
6 Limitations, future work, and conclusion
Limitations and future work.
We identify limitations, which provide motivation for future work.
  • Hyperparameters. While our configurations are surprisingly amenable to reliable scaling across many training and testing distributions without further tuning, there is a need to further develop scaling laws that incorporate hyperparameters.
  • Scaling up. Validating the trends in this paper for even larger runs is a valuable direction.
  • Scaling down. Additionally, actualizing predictable scaling with even cheaper runs is important to make this area of research more accessible, especially for downstream error prediction.
  • Failure cases. While we present preliminary analysis of when scaling is unreliable, future work should develop an analytic theory explaining when scaling breaks down.
  • Post-training. It is common to employ supervised fine-tuning and reinforcement learning after pre-training, which we do not consider. Quantifying to what degree over-training the base model provides benefits after post-training is an open area of research.
  • Individual downstream task prediction. While we find that averaging over many task error metrics can make for a predictable metric, per-task predictions are left to future work.
  • In-the-wild performance. Downstream task performance is a proxy for the in-the-wild user experience. Analyzing scaling trends in the context of this experience is timely.
  • Dataset curation. Our work only deals with existing training datasets. Exploring dataset curation for improved model scaling is another promising direction.
Conclusion.
We (i) show that the loss scaling behavior of models trained past compute-optimal, in the over-trained regime, is predictable and (ii) predict, via a proposed scaling law, the downstream average task performance of more expensive runs using smaller-scale proxies. We hope our work will inspire others to further examine the relationship between model training and downstream generalization. We also hope our testbed will make scaling research more accessible to researchers and practitioners alike.
Acknowledgements
SYG is supported by an NSF Graduate Research Fellowship, GS by the Onassis Foundation - Scholarship ID: F ZS 056-1/2022-2023, and MN by the Federal Ministry of Education and Research of Germany under grant no. 01IS22094B WEST-AI. We thank Stability AI and Toyota Research Institute (TRI) for access to compute resources. This research has been supported by NSF Grants AF 1901292, CNS 2148141, Tripods CCF 1934932, IFML CCF 2019844, and research gifts by Western Digital, Amazon, WNCG IAP, UT Austin Machine Learning Lab (MLL), Cisco, and the Stanly P. Finch Centennial Professorship in Engineering. We also thank Kushal Arora, Alper Canberk, Mia Chiquier, Sachit Menon, Chuer Pan, Purva Tendulkar, and Mandi Zhao for valuable feedback.
Appendix A Contributions
Names ordered alphabetically.
Planning.
Samir Yitzhak Gadre
Model training and experiment babysitting.
Achal Dave (notably, the 1.4B parameter, 900B token run), Samir Yitzhak Gadre
Dataloading.
Georgios Smyrnis
Training tokens.
Achal Dave, Alex Fang, Samir Yitzhak Gadre, Suchin Gururangan, Jeffrey Li, Vaishaal Shankar (lead), Mitchell Wortsman
Evaluation tokens.
Achal Dave, Samir Yitzhak Gadre, Reinhard Heckel, Vaishaal Shankar (lead), Rulin Shao
Loss/perplexity evaluation.
Samir Yitzhak Gadre
Downstream evaluation.
Vaishaal Shankar
Project-specific infrastructure, plots, and analysis.
Samir Yitzhak Gadre
OpenLM [37] open-source infrastructure.
Achal Dave (core contributor), Alex Fang, Samir Yitzhak Gadre (core contributor), Suchin Gururangan (core contributor), Jenia Jitsev, Sedrick Keh, Jeffrey Li, Jean Mercat, Marianna Nezhurina, Vaishaal Shankar (core contributor), Georgios Smyrnis (core contributor), Igor Vasiljevic, Mitchell Wortsman (core contributor), Rui Xin
Theory.
Yair Carmon (original idea that “parallel lines” should show up in scaling plots), Samir Yitzhak Gadre (various derivations, empirical verification, related validation loss to average top-1 error as in Equation (5)), Reinhard Heckel (derived a scaling form based on Chinchilla Approach 3 [43], which appears in Equation (4)), Mitchell Wortsman (provided intuition about irreducible loss and why it is critical), Niklas Muennighoff (derived a scaling form based on Chinchilla Approach 3, similar to Equation (4)).
Writing.
Yair Carmon, Achal Dave, Reinhard Heckel, Samir Yitzhak Gadre (lead), Niklas Muennighoff, Ludwig Schmidt
Compute.
Achal Dave, Jenia Jitsev, Thomas Kollar, Ludwig Schmidt, Vaishaal Shankar
Advising.
Yair Carmon (co-lead), Achal Dave (co-lead), Alexandros G. Dimakis, Reinhard Heckel (co-lead), Gabriel Ilharco, Jenia Jitsev, Thomas Kollar, Niklas Muennighoff (co-lead), Ludwig Schmidt (co-lead), Shuran Song
Appendix B Scaling-law derivations
We first show that reparameterizing Equation (3) in terms of the compute $C$ and token multiplier $M$ for $\alpha=\beta$ yields Equation (4). Combining $C=6ND$ and $M=D/N$ yields $N=\sqrt{C/(6M)}$ and $D=\sqrt{CM/6}$. Inserting these into Equation (3) yields,
$L(C,M)$ $=E+A\left(\frac{C}{6M}\right)^{-\frac{\alpha}{2}}+B\left(\frac{CM}{6}\right)^{-\frac{\alpha}{2}},$
$=E+\left(A\left(\frac{1}{6}\right)^{-\frac{\alpha}{2}}M^{\frac{\alpha}{2}}+B\left(\frac{1}{6}\right)^{-\frac{\alpha}{2}}M^{-\frac{\alpha}{2}}\right)C^{-\frac{\alpha}{2}}.$
This is equal to Equation (4), making the substitutions $\alpha_{C}=\alpha/2$, $a=A(1/6)^{-\alpha_{C}}$, $b=B(1/6)^{-\alpha_{C}}$, as noted in the main body.
Relation to compute-optimal training.
Recall that we made the assumption $\alpha=\beta$, which implies equal scaling of parameters and tokens to realize compute-optimal models. While this assumption is empirically justified [43], even if $\alpha\neq\beta$, we get a parameterization that implies the power law exponent in Equation (4) remains constant with over-training, while the power law scalar changes.
To find a compute-optimal training setting, Hoffmann et al. [43] propose to minimize the right-hand side of Equation (3) subject to the compute constraint $C=6ND$. This yields, $N^{\ast}=\gamma^{\frac{1}{\alpha+\beta}}(C/6)^{\frac{\beta}{\alpha+\beta}}$ and $D^{\ast}=\gamma^{-\frac{1}{\alpha+\beta}}(C/6)^{\frac{\alpha}{\alpha+\beta}}$, where $\gamma=\frac{\alpha A}{\beta B}$, for notational convenience. The associated risk is,
$L(N^{\ast},D^{\ast})=E+\left(A\gamma^{\frac{-\alpha}{\beta+\alpha}}+B\gamma^{\frac{\beta}{\beta+\alpha}}\right)\left(\frac{C}{6}\right)^{-\frac{\alpha\beta}{\alpha+\beta}}.$
We now deviate from compute-optimal training by modifying the model size and tokens by multiplication with a constant $\sqrt{m}$, according to
$N_{m}=\frac{1}{\sqrt{m}}N^{\ast},\quad D_{m}=\sqrt{m}D^{\ast}.$ (7)
This modification keeps the compute constant (i.e., $6N_{m}D_{m}=6N^{\ast}D^{\ast}$). The risk, then, becomes
$L(f_{N_{m},D_{m}})=E+\left(m^{\frac{\alpha}{2}}A\gamma^{\frac{-\alpha}{\beta+\alpha}}+m^{-\frac{\beta}{2}}B\gamma^{\frac{\beta}{\beta+\alpha}}\right)C^{-\frac{\alpha\beta}{\alpha+\beta}}.$ (8)
We again expect the same power law exponent and changing power law scalar. Note that $m$ in Equation (8) is similar to $M$ in Equation (4). Specifically, $m$ is a multiple of the Chinchilla-optimal token multiplier $M^{\ast}=D^{\ast}/N^{\ast}$, which is no longer fixed as a compute budget changes for $\alpha\neq\beta$.
Appendix C Additional grid search details
Grid search configuration selection.
Recall in Section 3.2, we run a grid search over many configurations. We present the architectures we sweep over in Table 5.
Table 5: Topologies for our grid searches. We consider 130 architectures for our grid search. After sweeping over batch size and warmup, we get a total of 435 configurations. For a complete list of hyperparameter configurations, please see: https://github.com/mlfoundations/scaling
$n_{layers}$ $n_{heads}$ $d_{model}$ Number of parameters [B]
4 4 96 0.010
4 12 96 0.010
12 12 96 0.011
12 4 96 0.011
8 4 96 0.011
16 4 96 0.011
16 12 96 0.011
8 12 96 0.011
24 4 96 0.012
24 12 96 0.012
4 4 192 0.021
4 8 192 0.021
4 12 192 0.021
8 8 192 0.023
8 4 192 0.023
8 12 192 0.023
12 4 192 0.025
12 8 192 0.025
12 12 192 0.025
16 4 192 0.026
16 8 192 0.026
16 12 192 0.026
24 8 192 0.030
24 4 192 0.030
24 12 192 0.030
4 12 288 0.033
4 4 288 0.033
8 12 288 0.037
8 4 288 0.037
4 4 320 0.038
4 8 320 0.038
12 12 288 0.041
12 4 288 0.041
8 8 320 0.043
8 4 320 0.043
16 4 288 0.045
16 12 288 0.045
12 4 320 0.049
12 8 320 0.049
24 4 288 0.053
24 12 288 0.053
16 8 320 0.055
16 4 320 0.055
4 12 488 0.062
4 4 512 0.065
4 16 512 0.065
4 8 512 0.065
24 8 320 0.066
24 4 320 0.066
4 4 576 0.074
4 8 576 0.074
4 12 576 0.074
8 12 488 0.075
8 4 512 0.079
8 8 512 0.079
8 16 512 0.079
4 4 640 0.085
4 16 640 0.085
4 8 640 0.085
12 12 488 0.087
8 4 576 0.090
8 12 576 0.090
8 8 576 0.090
12 16 512 0.093
12 8 512 0.093
$n_{layers}$ $n_{heads}$ $d_{model}$ Number of parameters [B]
12 4 512 0.093
16 12 488 0.100
8 16 640 0.105
8 4 640 0.105
8 8 640 0.105
12 8 576 0.106
16 16 512 0.106
4 4 768 0.106
12 12 576 0.106
16 8 512 0.106
4 8 768 0.106
12 4 576 0.106
4 16 768 0.106
16 4 512 0.106
4 12 768 0.106
16 12 576 0.122
16 4 576 0.122
16 8 576 0.122
12 4 640 0.126
24 12 488 0.126
12 16 640 0.126
12 8 640 0.126
24 8 512 0.133
24 4 512 0.133
24 16 512 0.133
8 8 768 0.134
8 16 768 0.134
8 4 768 0.134
8 12 768 0.134
16 16 640 0.146
16 8 640 0.146
16 4 640 0.146
24 8 576 0.154
24 4 576 0.154
24 12 576 0.154
4 8 1024 0.155
4 16 1024 0.155
4 4 1024 0.155
12 8 768 0.162
12 4 768 0.162
12 12 768 0.162
12 16 768 0.162
24 16 640 0.186
24 8 640 0.186
24 4 640 0.186
16 16 768 0.191
16 4 768 0.191
16 8 768 0.191
16 12 768 0.191
8 8 1024 0.206
8 4 1024 0.206
8 16 1024 0.206
24 8 768 0.247
24 12 768 0.247
24 4 768 0.247
24 16 768 0.247
12 8 1024 0.257
12 4 1024 0.257
12 16 1024 0.257
16 8 1024 0.309
16 4 1024 0.309
16 16 1024 0.309
24 16 1024 0.412
24 8 1024 0.412
24 4 1024 0.412
Table 5: Topologies for our grid searches. We consider 130 architectures for our grid search. After sweeping over batch size and warmup, we get a total of 435 configurations. For a complete list of hyperparameter configurations, please see: https://github.com/mlfoundations/scaling
Appendix D Evaluation dataset details
All 46 downstream evaluations are based on MosaicML’s LLM-foundry evaluation suite [67]. We specifically consider the datasets given in Table 6.
Downstream task LLM-foundry category Evaluation type Shots Samples Chance accuracy
AGIEval LSAT AR [129, 128, 116] symbolic problem solving multiple choice 3 230 0.25
AGIEval LSAT LR [129, 128, 116] reading comprehension multiple choice 3 510 0.25
AGIEval LSAT RC [129, 128, 116] reading comprehension multiple choice 3 268 0.25
AGIEval SAT English [129] reading comprehension multiple choice 3 206 0.25
ARC-Challenge [21] world knowledge multiple choice 10 2376 0.25
ARC-Easy [21] world knowledge multiple choice 10 2376 0.25
BBQ [77] safety multiple choice 3 58492 0.50
BIG-bench: CS algorithms [11] symbolic problem solving language modeling 10 1320 0.00
BIG-bench: Conceptual combinations [11] language understanding multiple choice 10 103 0.25
BIG-bench: Conlang translation [11] language understanding language modeling 0 164 0.00
BIG-bench: Dyck languages [11] symbolic problem solving language modeling 10 1000 0.00
BIG-bench: Elementary math QA [11] symbolic problem solving multiple choice 10 38160 0.25
BIG-bench: Language identification [11] language understanding multiple choice 10 10000 0.25
BIG-bench: Logical deduction [11] symbolic problem solving multiple choice 10 1500 0.25
BIG-bench: Misconceptions [11] world knowledge multiple choice 10 219 0.50
BIG-bench: Novel Concepts [11] commonsense reasoning multiple choice 10 32 0.25
BIG-bench: Operators [11] symbolic problem solving language modeling 10 210 0.00
BIG-bench: QA WikiData [11] world knowledge language modeling 10 20321 0.00
BIG-bench: Repeat copy logic [11] symbolic problem solving language modeling 10 32 0.00
BIG-bench: Strange stories [11] commonsense reasoning multiple choice 10 174 0.50
BIG-bench: Strategy QA [11] commonsense reasoning multiple choice 10 2289 0.50
BIG-bench: Understanding fables [11] reading comprehension multiple choice 10 189 0.25
BoolQ [19] reading comprehension multiple choice 10 3270 0.50
COPA [90] commonsense reasoning multiple choice 0 100 0.50
CoQA [89] reading comprehension language modeling 0 7983 0.00
Commonsense QA [105] commonsense reasoning multiple choice 10 1221 0.25
Enterprise PII classification [79] safety multiple choice 10 3395 0.50
HellaSwag (10-shot) [123] language understanding multiple choice 10 10042 0.25
HellaSwag (zero-shot) [123] language understanding multiple choice 0 10042 0.25
Jeopardy [67] world knowledge language modeling 10 2117 0.00
LAMBADA [75] language understanding language modeling 0 5153 0.00
LogiQA [56] symbolic problem solving multiple choice 10 651 0.25
MMLU (5-shot) [38] world knowledge multiple choice 5 14042 0.25
MMLU (zero-shot) [38] world knowledge multiple choice 0 14042 0.25
MathQA [5] symbolic problem solving multiple choice 10 2983 0.25
OpenBook QA [66] commonsense reasoning multiple choice 0 500 0.25
PIQA [12] commonsense reasoning multiple choice 10 1838 0.50
PubMed QA Labeled [48] reading comprehension language modeling 10 1000 0.00
SIQA [95] commonsense reasoning multiple choice 10 1954 0.50
SQuAD [88] reading comprehension language modeling 10 10570 0.00
Simple Arithmetic: NoSpaces [67] symbolic problem solving language modeling 10 1000 0.00
Simple Arithmetic: WithSpaces [67] symbolic problem solving language modeling 10 1000 0.00
WinoGender MC: Female [92] safety multiple choice 10 60 0.50
WinoGender MC: Male [92] safety multiple choice 10 60 0.50
WinoGrande [93] language understanding schema 0 1267 0.50
WinoGrand [53] language understanding schema 0 273 0.50
Table 6: 46 downstream tasks. All downstream tasks considered in this work, evaluated via LLM-foundry [67]. For more information on each dataset and specifics about the LLM-foundry category and evaluation type, please see: https://www.mosaicml.com/llm-evaluation.
Appendix E Additional results
Over-performing grid search models experience more optimization steps.
As mentioned in Section 3.2 and Figure 4, we notice that models between 0.011B to 0.079B (i.e., $5.2\times 10^{16}$ to $5.2\times 10^{17}$ FLOPs trained near compute-optimal) over-perform compared to the trend established by other models in our initial grid searches. This results in a bump in the scaling plot. While we choose to exclude this range of models for our scaling study, we additionally investigate this phenomenon. In Figure 6 we color grid search configurations by the number of optimization steps (i.e., number of tokens seen divided by batch size divided by sequence length). For context, Figure 1 (left) in Kaplan et al. [49] also shows a bump; however, there the performance is worse than the general trend instead of better as in our work. We leave understanding more fully the interactions between hyperparameters, scaling, and performance to future work.
Figure 6: Understanding over-performing models in our grid search. (left) Models trained with $5.2\times 10^{16}$ to $5.2\times 10^{17}$ FLOPs over-perform relative to their neighbors. In looking at the number of optimization steps, we notice that the over-performing models experience more optimization steps than their x-axis neighbors. We hypothesize that the number of optimization steps is important, especially for smaller models, when trying to find models that lie along a trend. (right) A view of the same phenomenon, specifically on the efficient frontier.
Scaling is largely predictable in-distribution (ID).
Prior work focuses on understanding scaling using ID loss, often using training loss directly [49, 43]. Hence, we also consider Paloma [63] loss evaluation sets, which are designed to probe performance in specific domains. We use Paloma’s C4 [86, 25], RedPajama [110], and Falcon-RefinedWeb [80] splits to probe for ID validation loss. As seen in Figure 7, relative error is mostly low. Relative error is largest for the $N=1.4\text{B},M=640$ RedPajama run at 15.4%. Examining this case specifically, we find that the model performs better than the scaling law prediction. We hypothesize that as a model sees more tokens there is an increased likelihood of near-duplicate sequences ID, resulting in performance that is better than predicted.
Figure 7: In-distribution (ID) settings. Boxes highlighted in yellow correspond to data points used to fit Equation (4). Relative error is generally low across interpolation and extrapolation regimes. Relative error is largest for the RedPajama $N=1.4\text{B},M=640$ prediction at 15.4%. In this case, we find that our scaling law predicts the model should perform worse than it does in practice.
Relative error is stable across many choices of downstream evaluation suites.
To understand how sensitive our investigation is to our choices of the evaluation set, we consider several other options as seen in Figure 8. We find that our prediction errors are fairly (i) low and (2) consistent for many choices of downstream evaluation sets.
Figure 8: Downstream evaluation set ablation for 6.9B parameter, 138B token runs. Recall that in addition to the 46 task evaluation suite, we consider a 17 task evaluation suite created by including only test sets where any 0.154B model we trained (for any token multiplier and training dataset) gets $t=10$ percentage points above random chance. Here we wish to understand how sensitive our results are to this choice $t$. (left) We see that relative prediction error is fairly low before a threshold of $t=35$. When too many tasks are excluded (i.e., $t>40$ relative error spikes. (right) A parallel view, showing how many tasks are removed as $t$ increases. 40 out of the 46 tasks can be removed and relative error is still fairly stable.
Scaling law fit Train set MMLU [38] OpenBook QA [66] HellaSwag [123] 17 eval 46 eval
Table 2 C4 [86, 25] 2.82% 16.80% 79.58% 0.14% 0.07%
Table 2 w/o 1.4B C4 [86, 25] 1.86% 96.16% 61.79% 0.42% 0.76%
Table 2 RedPajama [110] 0.12% 8.44% 25.73% 0.05% 2.10%
Table 2 w/o 1.4B RedPajama [110] 1.07% 7.56% 30.98% 10.64% 7.68%
Table 2 RefinedWeb [80] 0.77% 1.92% 81.96% 2.94% 2.76%
Table 2 w/o 1.4B RefinedWeb [80] 2.29% 6.79% 6.52% 15.79% 8.57%
Table 7: Downstream relative prediction error at 6.9B, 138B tokens, with and without the 1.4B data point. Recall in Table 2, we introduce a $N=1.4$B, $M=20$ run to get better downstream error predictions. Here we compare, prediction errors with and without this model for fitting the scaling law. Note that without the model (i.e., rows with “w/o 1.4B”) average top-1 predictions over a 17 task subset and over the full 46 task suite are less accurate.
Scaling can break down when under-training.
Figure 9: Scaling with small token multipliers. For smaller multipliers (e.g., $M=5$ in cyan), scaling does not follow the same trend as that of larger multipliers. Additionally, many token multiplies (e.g., $M\in\{10,20,40,80\}$) garner points close to the compute-optimal frontier.
We find that when a token multiple is too small (i.e., under-training regime), scaling appears unreliable. In Figure 9 we see for $M=5$ the scaling trend is different. We hypothesize that tuning hyperparameters (e.g., warmup, batch size) directly for smaller multipliers may help mitigate the breakdown in predictability.
Scaling can be unpredictable out-of-distribution (OOD).
Our main result shows reliable C4 eval loss predictions with models trained on RedPajama, which is an OOD evaluation setting. However, both C4 and RedPajama both contain tokens sourced from CommonCrawl.
To further probe OOD performance, we measure the relative error of scaling laws fit to models trained on C4 and evaluated on Paloma’s 100 programming languages [63], Paloma’s Penn Tree Bank (PTB) split [64], and a German version of C4 [25]. Recall that the C4 training set we use has been filtered for English text. Hence we expect (i) the proportion of code is minimal, (ii) the “<unk>” substrings in PTB raw text do not appear frequently, and (iii) German is not prevalent. We notice that extrapolation relative error tends to be high for large $M,N$ on programming languages and PTB (Figure 10 (left, center)). In contrast, for German C4, relative error is still low across the extrapolation range with a maximum relative error of 7.6% at the $N=$1.4B, $M=80$ scale (Figure 10 (right)). We hypothesize that further modifications to scaling laws are necessary to predict when scaling should be reliable as a function of the training and evaluation distributions.
Figure 10: Out-of-distribution (OOD) settings. Boxes highlighted in yellow correspond to data points used to fit Equation (4). Recall that the C4 training set is English-filtered. Relative error can spike, suggesting unreliable scaling, for (left) programming languages and (center) Penn Tree Bank, which contains many frequently occurring, uncommon substrings. However, scaling is relatively reliable when evaluating on (right) German. These results motivate future studies of OOD conditions that affect scaling in the over-trained regime.
Small-scale experiments can predict average downstream top-1 error.
Figure 11: Relative error on average top-1 predictions (46 task split). Boxes highlighted in yellow correspond to data points used to fit Equation (5). Using our fits, we accurately predict downstream average top-1 error across interpolation and extrapolation regimes. This result supports that (i) chaining a scaling law and our proposed exponential decay function is a valid procedure and (ii) average top-1 error can be highly predictable.
To verify that chaining Equations (4) and (5) is effective in practice, we collect C4 eval loss and downstream error pairs for the configurations in Table 2. In Figure 11, we look at relative error for our scaling predictions in the context of Average top-1 error over our 46 evals and over our 17 evals in Figure 12. We again notice reliable scaling in interpolation and extrapolation regimes, suggesting the validity of our procedure to predict downstream average top-1 error.
Figure 12: Relative error on average top-1 predictions (17 task split). Boxes highlighted in yellow correspond to data points used to fit Equation (5). Using our fits, we accurately predict downstream average top-1 error across interpolation and extrapolation regimes. This result supports that (i) chaining a scaling law and our proposed exponential decay function is a valid procedure and (ii) average top-1 error can be highly predictable.
Loss evaluation ablations for downstream trends.
Figure 13 presents the correlation between downstream error vs. loss evaluated on different validation sets (C4, RedPajama, and RefinedWeb). Regardless of the validation set (x-axis), models follow the exponential decay relationship given in Equation (5), suggesting the choice of validation loss is not critical for the appearance of this phenomenon.
Figure 13: Correlation between average top-1 error and evaluation loss. We observe that regardless of evaluation loss distribution (x-axis), models tend to follow Equation (5). This suggests that there can be several reasonable choices for the validation loss distribution. Additionally, ID models trained on C4 and evaluated on a C4 validation set, perform best in terms of loss, but these gains don’t necessarily translate to lower error downstream (e.g., (left column)). This suggests the need to fit Equation (5) per dataset and also suggests comparing models trained on different data distributions with a single loss evaluation can be misleading in terms of downstream performance.
Investing more compute in a scaling law makes it more predictive.
Figure 14: Trade-offs between scaling law for loss fitting considerations and reliability. Each red circle represents a scaling law fit to Equation (4) with as many as 29 models trained on RedPajama. Specifically, a grid formed by $N\in\{0.011\text{B},0.079\text{B},0.154\text{B},0.411\text{B}\},M\in\{5,10,20,40,80,160,320\}$ gives 28 models and a $N=1.4B,M=20$ run gives the last model. We sort models by training FLOPs in increasing order and sample models uniformly from index windows $[1,2,...,n]$ for $n\in[5,6,..,29]$ to fit Equation (4). The blue star represents the default configuration presented in Table 2. The prediction target is a $N=1.4B,M=640\text{ }(D=900\text{B})$ model. As the amount of compute (left) and the number of points (right) used to fit the scaling law increases, relative error trends downwards. Our default configuration keeps compute and number of points low, while still providing low prediction error compared to the trend.
Figure 15: Compute vs. relative error for the 1.4B, 900B token RedPajama run. (left) The compute necessary to accurately predict loss is less than that needed to accurately predict (right) average downstream error. This claim is supported by the fact that the slope of the trend for loss is steeper than for top-1 error. These findings corroborate Figure 16.
Figure 16: Compute vs. relative error for the 6.9B, 138B token RedPajama run. (left) The compute necessary to accurately predict loss is less than that needed to accurately predict (right) average downstream error. This claim is supported by the fact that the slope of the trend for loss is steeper than for top-1 error. These findings corroborate Figure 15.
Thus far we have looked at standard configurations from Table 2 to construct our scaling laws, mainly to demonstrate extrapolation to larger $N,M$. However, for practitioners, the main constraint is often training compute. Hence, we wish to understand the trade-offs between the amount of compute invested in creating a scaling law and the relative error of the resulting law in the over-trained regime. In Figure 14 (left), we see that as one increases the amount of compute, it is possible to get better fits with lower relative error. In Figure 14 (right), we see a similar trend as one increases the number of data points used to fit a scaling law. Blue stars indicate the configurations from Table 2, which provide accurate predictions relative to the general trends—hinting at their usefulness for our investigation. In Figures 1516 we repeat the compute analysis comparing trade-offs for loss prediction and error prediction reliability. We find that less compute is generally necessary to construct a loss scaling law that achieves the same relative error as that of an error prediction scaling law.
Figure 17: Scaling exponent vs. token multiplier. In Figure 2 we notice roughly parallel lines (i.e., roughly constant scaling exponent $\alpha_{C}$) in the $\log$-$\log$ plot of loss vs. compute, even as the token multiplier $M$ changes. Here we plot $\alpha_{C}$ vs. $M$ directly, where the shaded region gives a 95% bootstrap confidence interval for the trend. This view supports that $\alpha_{C}$ is relatively constant.
Figure 18: Downstream top-1 error vs. C4 eval loss for each of the 46 downstream evals. Here we plot models from our testbed for each scatter plot. We see that some individual evaluations, like ARC-Easy, follow exponential decay. Others, like BIG-bench: CS algorithms, show step function behavior. Still others, like MathQA, hover around random chance.
Appendix F Additional related work
Language modeling.
Language models can be grouped into encoder-only [24, 51, 57, 94, 20], encoder-decoder [54, 87], and decoder-only architectures [83, 111, 112, 108, 47, 36, 72, 7, 109, 26, 62, 97, 119, 4, 55, 61, 32]. Most current implementations are based on the transformer [114]. However, there has been a recent resurgence in scaling language models based on non-transformer architectures [81, 34, 35, 33]. Further, there has been substantial work on adapting pre-trained language models to better follow instructions [117, 18, 68, 59, 69, 130, 85, 27, 113, 101, 71]. However, following prior work [43, 70] and given their overall prevalence, we limit ourselves to GPT-style, decoder-only transformers that have solely been pre-trained.
Scaling laws.
Kaplan et al. [49] investigate scaling trends in GPT language models. Bahri et al. [9] investigate different scaling regimes theoretically, and Sharma & Kaplan [99] relate scaling coefficients to data manifold dimensions. Tay et al. [106, 107] elucidate the connection between model architecture and scaling trends, while Hernandez et al. [40], Tay et al. [106] develop scaling laws for transfer learning. Ivgi et al. [46] also consider transfer learning scaling laws and highlight the importance of hyperparameter selection in the low-compute regime. Ghorbani et al. [30], Gordon et al. [31], Bansal et al. [10] develop scaling laws for neural machine translation. Caballero et al. [15] propose a scaling law functional form, which they demonstrate is predictive in several domains.
Scaling beyond language modeling.
There is a large body of work on scaling neural networks beyond language modeling, for example in computer vision [58, 124, 103, 1, 2], multimodal learning [39, 16, 28], and image reconstruction [50].